The following table shows the actual demand observed over the last 4 years:                                                                                                                                                                                                                                       

Year

1

2

3

4

Demand

8

10

6

9


Using exponential smoothing with alpha α=0.30 and a forecast for year 1 of 7.0, provide the forecast from periods 2 through 5 (round your responses to one decimal place).

Year

1

2

3

4

5

Forecast (ES)

7.0






Similarly calculate the forecast for each of the years from 3 to 5 using the naive approach (enter your responses as whole numbers).

Year

2

3

4

5

Forecast (NA)

8
Answer #1

Using the exponential smoothing method

F(2) = 0.3*Actual(1) + (1-0.3)*F(1)

= 0.3*8 + 0.7*7

= 7.3

F(3) = 0.3*10 + 0.7*7.3

= 8.1

F(4) = 0.3*6 + 0.7*8.1

= 7.5

F(5) = 0.3*9 + 0.7*7.5

= 8

Using the naive method

F(2) = Actual(1)

= 8

F(3) = Actual(2)

= 10

F(4) = 6

F(5) = 9

Please provide up votes, if you are satisfied with the response. Comment in case of any issue regarding the same.